# GNSS-inertial state initialization by distance residuals

### v1

@@ V-A Experimental setup
All the experiments were run on an Intel Core equipped with an i7-11700K CPU at 3.6 GHz and 64 GB memory. Our implementation leverages the open-source PyPose library [32], which provides Lie group operations and optimization routines tailored for robotics and state estimation. The datasets used in the experiments were selected from public datasets that provide IMU, GPS measurements and ground-truth data. For all experiments, we report the Root Mean Square Error of the Absolute Trajectory Error (ATE RMSE $\downarrow$) to assess the position error. We chose the public datasets Euroc and GVINS, as they are typically used in relevant recent works involving GNSS and inertial data [33].
V-B EuRoc dataset
The EuRoc MAV dataset [34] provides high-quality synchronized data from IMU, stereo cameras, and motion capture ground-truth in indoor environments. Since GNSS data is not available, we follow the methodology in [33] and simulate noisy GNSS measurements by adding zero-mean Gaussian noise with a standard deviation of $\sigma_{\text{g}}=0.2~{}m$ to the ground-truth positions. The aim of this experiment is to assess the impact of optimizing the rigid transformation matrix $\mathbf{T}_{W}^{B}$ between the GPS and IMU frames on the overall performance of the system. A series of experiments are conducted in which this transformation is optimized, but only after a predefined minimum number of GPS measurements have been accumulated. This threshold is varied across experiments and is represented on the x-axis. Fig. 6 shows the aforementioned results. Performance is evaluated in terms of position error. For this sequences, the full set of 100 GNSS measurements corresponds to a trajectory of approximately $20~{}m$.
Figure 6: RMSE values for different EuRoc sequences. x-axis represent the number of GPS measurements accumulated before including global residuals in the optimization.
The experimental results reveal a consistent trend: deferring the inclusion of the rigid-body transformation $\mathbf{T}_{W}^{B}$ in the optimization, until a minimum number of GPS measurements have been accumulated, leads to substantially lower position errors. This delay enables a more robust and accurate estimation of the transformation parameters, effectively mitigating the influence of early measurement noise and initial uncertainty. In contrast, the worst-case scenario, in which the transformation is entirely excluded from the optimization, results in significantly degraded performance. This highlights the importance of accurately calibrating the spatial relationship between the inertial and global frames for reliable state estimation.
V-C GVINS dataset
The GVINS dataset [35] features three real outdoor sequences recorded with a multi-sensor setup mounted on a UAV, IMU, and high-rate RTK-GNSS. The sequences contained are Sports field, Indoor-outdoor and Urban driving. In our experiments, we preprocess the data by extracting the IMU and GPS measurements and aligning the latter into the UTM coordinate system. The objective of this experiment is to evaluate the proposed optimization framework in more realistic, large-scale scenarios and to analyze the influence of delayed activation of global residuals on overall estimation accuracy. To this end, we vary the activation index for including global GPS residuals in the optimization and report the resulting trajectory errors in Fig. 7. The full set of 100 GPS measurements corresponds to a trajectory length of approximately $13~{}{m}$ in the Sports field and Indoor-outdoor sequences, and about $35~{}{m}$ for the Urban driving sequence.
Figure 7: RMSE values for different GVINS sequences. x-axis represent the number of GPS measurements accumulated before including global residuals in the optimization.
As in the previous experiment, a similar pattern is observed in the GVINS dataset. Across the three representative sequences, deferring the optimization of the transformation $\mathbf{T}_{W}^{B}$ until a sufficient number of GPS measurements have been collected consistently results in lower ATE RMSE. The benefit of delayed activation is especially pronounced in the Urban_driving sequence, where premature optimization leads to large estimation errors. Conversely, excessive delay in incorporating global constraints also results in increased trajectory error, as relevant global information is excluded for too long. These findings highlight the importance of introducing global residuals only when the estimation problem is sufficiently well-conditioned. Identifying this optimal activation point is therefore crucial, as both premature and overly delayed inclusion can negatively impact final estimation accuracy.
To validate this strategy more systematically, we next present experiments using the activation criterion introduced previously in Section IV-C.
V-D Triggering Criterion experiments
In the previous section, we introduced a criterion to determine the appropriate instant at which to incorporate global position residuals into the optimization. Fig. 8 shows the evolution of the relative change $\Delta\rho_{k}$ for two EuRoC sequences. When the curve drops below the observability threshold $\Delta\rho_{\mathrm{th}}=10^{-2}$, the system is considered sufficiently constrained to support the inclusion of global position residuals in the estimation.
@@ V-D Triggering Criterion experiments
Figure 8: Example of our activation criterion for two EuRoc sequences. a) On the MH05 sequence, the criterion results in $k^{*}=9$ GPS measurements; b) on the MH03 sequence, it suggests $k^{*}=15$ GPS measurements. The dashed blue line represents the observability threshold $\Delta\rho_{{th}}$.
@@
EuRoC MH_01_easy 11 0.058 0.057
MH_02_easy 12 0.051 0.049
MH_03_medium 15 0.051 0.040
MH_04_difficult 21 0.044 0.043
MH_05_difficult 9 0.049 0.050
V1_01_easy 20 0.048 0.046
V1_02_medium 10 0.041 0.040
V1_03_difficult 30 0.053 0.052
V2_01_easy 21 0.051 0.050
V2_02_medium 10 0.061 0.057
V2_03_difficult 12 0.042 0.040
@@
TABLE I: $k^{*}-$index obtained by our criterion and position error comparison for each sequence. ”Pos. Error (Full)” refers to ATE RMSE [$\downarrow~{}{m}$] when using GPS residuals from the beginning, while ”Pos. Error (from $k^{*}$)” shows the result when global residuals are introduced at $k^{*}$.
Table I reports the estimated activation index $k^{*}$ and compares the position error using two strategies: including GPS residuals from the beginning, or activating them at the recommended index $k^{*}$.
Our method achieves consistent improvements across both EuRoC and GVINS datasets. On average, this results in a reduction of trajectory error by $4\%$, with gains reaching up to $21\%$ in certain sequences such as MH_03_medium or Indoor-outdoor. These results validate the effectiveness of the proposed activation criterion, as it avoids premature inclusion of global residuals when the system is still underconstrained, thereby improving overall estimation accuracy.
@@ VI Conclusions
In this paper, we introduced a GNSS-inertial initialization method formulated through relative distance residuals. We also proposed a principled activation criterion that defers the inclusion of global residuals until the problem is sufficiently well-conditioned, as determined by an observability analysis. Experiments on the EuRoC and GVINS datasets confirm that this strategy yields accurate and robust trajectory estimates under a variety of operating conditions.

### v2

@@ V-A Experimental setup
All experiments were run on an Intel Core i7-11700K CPU (3.6 GHz, 64 GB RAM) using the open-source PyPose library [28] for Lie group operations and optimization. We assess tracking performance using the Root Mean Square of the Absolute Trajectory Error (ATE RMSE $\downarrow$) Besides trajectory accuracy, we report RMSE of gyro bias estimates and heading (yaw) error after initialization, following [29]. The EuRoC and GVINS datasets were selected as standard benchmarks in GNSS–inertial fusion [30]. We additionally evaluate on the MARS LVIG dataset, which contains large-scale outdoor trajectories. Sequences collected including manual flights at a low-altitude ($20~\mathrm{m}$) were used in our experiments. Unlike EuRoC and GVINS, it features long GNSS outages and stronger dynamics, providing a more challenging benchmark for testing the robustness of our initialization.
V-B Synthetic experiments
We conduct two complementary analyses on synthetic GNSS data, designed to isolate and better understand the influence of different modeling choices within our optimization framework.
First, we evaluate two alternative formulations of GNSS residuals within the same nonlinear optimizer and under a fixed-size sliding window. In the absolute variant (A1), we introduce absolute position constraints at each epoch, defined as $r_{k}^{\mathrm{abs}}=\mathbf{p}^{\mathrm{GNSS}}_{k}-\mathbf{p}_{k}(\mathbf{x})$. In the inter-epoch variant (A2), we instead enforce consistency between consecutive GNSS baselines, with residuals $r_{k}^{\mathrm{base}}=(\mathbf{p}^{\mathrm{GNSS}}_{k}-\mathbf{p}^{\mathrm{GNSS}}_{k-1})-(\mathbf{p}_{k}(\mathbf{x})-\mathbf{p}_{k-1}(\mathbf{x}))$. Both use identical weights, robust loss, window length, and stopping criteria, ensuring a fair comparison. The evaluation is carried out on EuRoC by varying the GNSS noise level from $\sigma_{p}=0.2$ to $2.0\,\mathrm{m}$, repeating each configuration ten times for statistical significance.
$\sigma_{p}$ $\mathrm{m}$] A1 Pos. $\downarrow\mathrm{m}$] A2 Pos. $\downarrow\mathrm{m}$] $\Delta\%$ (A2-A1)
0.2 0.05 $\pm$ 0.01 0.05 $\pm$ 0.01 0.0
1.0 0.53 $\pm$ 0.03 0.51 $\pm$ 0.04 3.8-3.8
2.0 1.64 $\pm$ 0.08 1.44 $\pm$ 0.08 12.2-12.2
[…]
@@ V-B Synthetic experiments
Figure 6: Example of our activation criterion for two EuRoc sequences. a) On the MH05 sequence, the criterion results in $k^{*}=9$ GNSS measurements; b) on the MH03 sequence, it suggests $k^{*}=15$ GNSS measurements. The dashed blue line represents the observability threshold $\Delta\rho_{{th}}$.
As shown in Table I, the two strategies yield similar performance under low-noise conditions. However, as GNSS quality deteriorates, the inter-epoch formulation (A2) proves more robust, leading to lower trajectory errors while maintaining essentially unchanged heading and gyroscope-bias estimates. These results suggest that inter-epoch baseline residuals are better suited for GNSS-challenged regimes, a design choice further validated with real GNSS data in the following subsections.
Second, we assess the effect of delaying the activation of global residuals and optimizing the rigid transformation $\mathbf{T}_{WB}$ using: (i) EuRoC MAV [31], where GNSS is synthetically generated by adding zero-mean Gaussian noise with $\sigma_{\text{p}}=0.2~\text{m}$ to motion-capture ground truth following [30]; and (ii) GVINS [32], which provides real outdoor IMU and high-rate RTK-GNSS. For GVINS, we extract IMU/GNSS data and align positions to the UTM frame.
A central question in this setup is when to activate global position residuals. Incorporating them too early, while the system is weakly constrained or underconstrained, can degrade accuracy. Delaying their incorporation until sufficient GNSS measurements are available allows the optimizer to use them more effectively. To address this, we proposed an observability-based triggering criterion that determines the appropriate instant to introduce global constraints. Fig. 6 illustrates the evolution of the relative change $\Delta\rho_{k}$ for two EuRoC sequences. When the curve drops below the observability threshold $\Delta\rho_{\mathrm{th}}=10^{-2}$, the system is considered sufficiently constrained to support the inclusion of global residuals. For instance, the criterion suggests $k^{*}=9$ GNSS measurements for MH05 and $k^{*}=15$ for MH03.
Finally, Fig. 7 shows the impact of this strategy on position accuracy. The x-axis in the figure denotes the number of GNSS measurements accumulated before enabling global residuals in the optimization. The full set of 100 measurements corresponds to a trajectory of approximately $20~\text{m}$. Activating global residuals prematurely increases error, whereas deferring them according to the proposed criterion leads to significantly lower trajectory error. These results highlight the importance of delaying global constraints to mitigate the effects of early noise and uncertainty.
@@
EuRoC MH01 11 0.058 0.057
MH02 12 0.051 0.049
MH03 15 0.051 0.040
MH04 21 0.044 0.043
MH05 9 0.049 0.050
V101 20 0.048 0.046
V102 10 0.041 0.040
V103 30 0.053 0.052
V201 21 0.051 0.050
V202 10 0.061 0.057
V203 12 0.042 0.040
@@
TABLE II: $k^{*}$-index obtained by our observability criterion and position errors. ‘Pos. Error (Full)’ refers to ATE RMSE $\downarrow{\mathrm{m}}$] when using GNSS global residuals from the outset, while ‘Pos. Error (from $k^{*}$)’ shows results when global residuals are introduced at $k^{*}$. Note consistent improvements by delaying global residuals until $k^{*}$.
Figure 7: RMSE values for different EuRoc sequences. x-axis represent the number of GNSS measurements accumulated before including global residuals in the optimization.
Table II extends this analysis by reporting the activation index $k^{*}$ and the resulting errors across EuRoC and GVINS. Our method consistently achieves lower position errors than the baselines, with the largest gains in sequences such as MH_03 and Indoor-outdoor. On average, trajectory error is reduced by about $4\%$, reaching up to $20\%$ in specific cases. These results validate the observability-based activation strategy, showing that postponing the inclusion of global residuals until the system is sufficiently constrained improves both accuracy and robustness across diverse datasets.
Heading Err. $\downarrow\mathrm{deg}$] $\mathbf{b}^{g}$ Err. $\downarrow\mathrm{rad/s}$]
Sequence EKF RA-JAPE Ours EKF RA-JAPE Ours
airport1 1.64 1.61 1.59 0.048 0.044 0.040
airport2 1.73 1.70 1.68 0.052 0.048 0.044
airport3 1.79 1.78 1.75 0.052 0.044 0.045
island1 1.30 1.31 1.32 0.040 0.040 0.038
island2 1.41 1.40 1.39 0.040 0.040 0.036
island3 1.56 1.52 1.50 0.048 0.044 0.040
town1 1.57 1.54 1.52 0.044 0.044 0.040
town2 1.66 1.60 1.61 0.048 0.044 0.048
town3 1.74 1.73 1.70 0.048 0.046 0.045
TABLE III: MARS-LVIG data evaluation. Per-sequence heading $\mathrm{deg}$] and gyroscope-bias $\mathrm{rad/s}$] errors for EKF, RA-JAPE [10], and Ours. Best per row is bolded.
V-C MARS-LVIG experiments
We further validate our approach on the MARS-LVIG dataset [33], which provides multi-sensor UAV sequences with data collected across diverse large-scale outdoor scenes. We follow the same experimental setup, but in this case we report errors in heading and gyroscope bias, comparing our method against state-of-the-art techniques: EKF and RA-JAPE [10]. This evaluation allows us to assess not only the impact on position accuracy, but also the ability of the proposed activation strategy to improve attitude estimation and bias calibration under realistic field conditions. Across the nine UAV sequences in Table III, the largest benefits appear on gyroscope bias: improvements reach up to +16.7% versus EKF on airport1 and island3, and +10.0% versus RA-JAPE on island2. Averaged across sequences, bias error decreases by +11% versus EKF and +5% versus RA-JAPE, with best results in 7/9 sequences. For heading, the best-case gains are +4% versus EKF on island3 and +2% versus RA-JAPE on town3; averaged across sequences the gains are +3% and +1%, respectively, and we obtain the lowest heading in 6/9 sequences. A few reversals remain (e.g., island1/island2 for heading, town2 and airport3 for bias), which we attribute to more challenging GNSS geometry or intermittent availability. While small in magnitude, such reductions are meaningful: sub-degree yaw errors at initialization can accumulate into drift [34]. The improved heading and bias estimates therefore yield a more reliable initial state for GNSS–inertial fusion.
Overall, the trend indicates that inter-epoch baseline GNSS residuals with activation substantially improve bias estimation while maintaining competitive heading accuracy under realistic, large-scale flights.
@@ VI Conclusions
In this paper, we introduced a GNSS-inertial initialization method formulated through inter-epoch baseline residuals. We also proposed a principled activation criterion that defers the inclusion of global residuals until the problem is sufficiently well-conditioned, as determined by an observability analysis. Experiments on EuRoC, GVINS and MARS-LVIG datasets confirm that this strategy yields accurate and robust trajectory estimates under a variety of operating conditions. Furthermore, our approach generalizes effectively to realistic GNSS deployments, providing consistent improvements not only in position accuracy but also in heading and gyroscope-bias estimation. Taken together, these results validate the proposed method as a reliable initialization strategy across both controlled benchmarks and challenging real-world scenarios.